### v1

# Thermal Inertia Controls on Titan’s Surface Temperature and Planetary Boundary Layer Structure

Abstract
Understanding Titan’s planetary boundary layer (PBL)—the lowest region of the atmosphere influenced by surface conditions—remains challenging due to Titan’s thick atmosphere and limited observations. Previous modeling studies have also produced inconsistent estimates of surface temperature, a critical determinant of PBL behavior, often without clear explanations grounded in surface energy balance. In this study, we develop a theoretical framework and apply a three-dimensional dry general circulation model (GCM) to investigate how surface thermal inertia influences surface energy balance and temperature variability across diurnal and seasonal timescales. At diurnal timescales, lower thermal inertia surfaces exhibit larger temperature swings and enhanced sensible heat fluxes due to inefficient subsurface heat conduction. In contrast, at seasonal timescales, surface temperature variations show weak sensitivity to thermal inertia, as atmospheric damping tends to dominate over subsurface conduction. The PBL depth ranges from a few hundred meters to 1,000 [m] on diurnal timescales, while seasonal maxima reach 2,000–3,000 [m], supporting the interpretation from a previous study that the Huygens probe captured the two PBL structures. Simulated seasonal winds at the Huygens landing site successfully reproduce key observed features, including near-surface retrograde winds and meridional wind reversals within the lowest few kilometers, consistent with Titan’s cross-equatorial Hadley circulation. Simulations at the planned Dragonfly landing site predict shallower thermal PBLs with broadly similar wind patterns. This work establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability, and offers a unified explanation of Titan’s PBL behavior that provides improved guidance for future missions.
1 Introduction
The planetary boundary layer (PBL) is the lowest atmospheric layer on terrestrial worlds, and is characterized by significant fluxes of heat, moisture, and momentum between the surface and atmosphere, driven by turbulent motions. On Earth, the PBL undergoes substantial diurnal and seasonal variations due to surface heating and convection. These transitions between temperature extrema, whether diurnal or seasonal, generate distinct atmospheric stability regimes that influence critical weather and climate processes, including cloud formation, precipitation, local breezes, and thunderstorms (Stull, 1988; Garratt, 1992; Lee, 2016). The phenomenon of the PBL extends beyond Earth to other celestial bodies possessing both a solid surface and an atmosphere (Petrosyan et al., 2011; Lebonnois et al., 2013; Flasar et al., 2014). As such, studying the PBL of other planetary bodies provides valuable insights into general PBL processes, making it crucial for comparative planetology. Mars, for instance, exhibits a PBL similar to Earth’s, where convection plays a fundamental role (Hinson et al., 2010; Spiga et al., 2010).
Titan, Saturn’s largest satellite, presents unique geological landscapes with dry, vast linear dunes at the equator contrasting with lakes and seas in the polar regions (Lorenz et al., 2006; Aharonson et al., 2014; Hayes, 2016). These striking geographical contrasts make understanding Titan’s PBL essential for explaining the formation and evolution of such diverse landscapes. Due to Titan’s vast distance from the Sun and its thick, opaque atmosphere, average surface insolation is estimated at approximately 0.1 % of Earth’s, around 1.0 [Wm-2] (McKay et al., 1989, 1991; Tomasko et al., 2008). However, given Titan’s extended day (16 Earth days) and year (29.5 Earth years), the accumulated surface energy over these cycles may suffice to generate notable surface temperature variations and PBL responses. Observational evidence remains limited, especially for a diurnal cycle, with only one in-situ vertical thermal profile near the equatorial surface from the Huygens landing (Fulchignoni et al., 2005) and scattered radio occultation profiles across the moon (Flasar et al., 1981; Lindal et al., 1983; Samuelson et al., 1997; Courtin & Kim, 2002; Schinder et al., 2011, 2012, 2020). Consequently, modeling studies are vital to supplementing the limited observational data and revealing whether Titan’s PBL responds to surface heating through convective processes in ways analogous to Earth and Mars, both diurnally and seasonally.
Understanding how surface temperature responds to solar forcing is essential for interpreting variations in the PBL. Within the context of modeling surface temperature variations on terrestrial planets, thermal inertia, defined as [JmsK-1] (hereafter we will use thermal inertia units; TIU), serves as a crucial parameter for quantifying a surface’s thermal response to energy input (Price, 1977; Putzig et al., 2005; MacKenzie et al., 2019). Here, represents the density [kgm-3], denotes the specific heat capacity [JkgK-1], and is the thermal conductivity [WmK-1] of the surface. Materials with lower thermal inertia tend to exhibit larger temperature variations, reflecting their diminished capacity to store and conduct heat (Stull, 1988; Garratt, 1992).
Using a dry general circulation model (GCM) of Titan’s atmosphere coupled with a simplified subsurface scheme, Tokano (2005) reported diurnal and seasonal surface temperature variations of a few kelvins near the equator—close to the latitude of the Huygens landing at 10.3°S—assuming a uniform thermal inertia of [TIU], characteristic of porous ice regolith. Williams et al. (2012) employed a one-dimensional energy-balance model, considering the Huygens landing site with a fixed near-surface atmospheric temperature, to investigate surface temperature variations. Their results indicated smaller diurnal and seasonal surface temperature variations, with amplitudes less than 0.1 [K] under dry conditions and a thermal inertia of [TIU], contradicting the findings of Tokano (2005). Subsequently, Charnay & Lebonnois (2012) reported diurnal fluctuations around 0.3 [K] in surface temperatures at the Huygens landing site using a dry GCM with [TIU] (Lebonnois et al., 2012). They proposed that the Huygens measurement of a near-adiabatic layer reaching 300 [m] (Tokano et al., 2006) reflected ongoing convection during the probe’s descent (at approximately 10 am local time), while changes in the lapse rate at 800 [m] and 2,000 [m] corresponded to residual signatures of maximum diurnal PBL extension and seasonal Hadley circulation effects, respectively. More recently, MacKenzie et al. (2019) implemented a global thermal inertia map in a dry GCM (Lora et al., 2015), demonstrating that seasonal surface temperature patterns show limited sensitivity to local thermal inertia values.
Previous studies reveal significant inconsistencies in our understanding of Titan’s surface temperature and, consequently, PBL variations. Notably, they report differing magnitudes of diurnal surface temperature variations even when assuming the same or similar values of thermal inertia. Equally puzzling is that seasonal surface temperature variations show little to no sensitivity to thermal inertia within a model, contradicting the expectation that lower thermal inertia surfaces should experience larger temperature swings. The underlying causes of these discrepancies remain unclear. A central limitation of prior work is the absence of a systematic analysis linking surface temperature variations to the governing surface energy balance.
Building on these insights, this study develops a theoretical framework that links surface temperature variations to thermal inertia and surface energy balance, and conducts dry GCM simulations to investigate how varying thermal inertia influences surface temperature and PBL behavior on Titan across diurnal and seasonal timescales. Section 2 introduces the theoretical foundations governing surface temperature variability and outlines the modeling approach. Section 3 presents GCM simulation results for surface temperature variations and demonstrates how the theoretical framework explains them at both diurnal and seasonal timescales. Section 4 examines PBL structures from GCM simulations and their relationship to surface temperatures. Section 5 evaluates expected conditions at the Dragonfly landing site. Section 6 discusses the implications of the findings and and concludes by highlighting the distinct roles of atmospheric damping and subsurface heat conduction in regulating surface temperature and PBL structures—dominant at seasonal and diurnal timescales, respectively. Overall, this study provides a critical reassessment of the atmospheric profile measured by Huygens and offers new insights into Titan’s lower atmosphere, informing expectations for the upcoming Dragonfly mission.
2 Methods
2.1 Theory
Energy balance at the surface for a dry case (i.e., without latent heat flux) is given by:
where and [Wm-2] represent the net radiative flux into and the sensible heat flux out of the surface, respectively. The residual, [Wm-2], is the net energy flux at the surface available for subsurface heat conduction.
The evolution of surface temperature follows the subsurface heat diffusion equation using as the surface boundary condition:
where [JmK-1] collectively denotes volumetric heat capacity. Here, homogeneous subsurface properties are considered for simplicity and to compare with previous results. The net energy flux at the surface can be simplified as a single harmonic oscillator with linear atmospheric damping (e.g., Dickinson, 1988):
where [Wm-2] is the amplitude of the periodic flux, and [Wm-2K-1] is the atmospheric damping coefficient. Over land surfaces, as assumed in dry GCM configurations, an additional lower boundary condition must be applied to the subsurface; for example:
where [K] is some mean temperature in the deep subsurface, and [m] is the thermal skin depth corresponding to a forcing frequency of [s-1], whether it be diurnal or annual. The magnitude of surface temperature fluctuations is expressed as (see Appendix A for the full derivation):
where [Wm-2] denotes the amplitude of fluctuations in . The thermal inertia [TIU], combined with [s-0.5] represents damping of the energy ([Wm-2K-1]) received at the surface, due to subsurface heat conduction. Meanwhile, [Wm-2K-1], which is the magnitude of , accounts for damping due to direct energy exchange between the surface and the atmosphere.
For Titan, the temperature of the overlying atmosphere, [K], can be reasonably assumed to remain constant due to its long radiative timescale (Flasar et al., 1981, 2014). Notably, the linear atmospheric damping term in Eq. (4) takes the same form as the expression for sensible heat flux:
where [kgm-3] is the density of Titan’s lower atmosphere as measured by the Huygens probe, is the specific heat capacity of the N2-dominated atmosphere [JkgK-1], [dimensionless] is the surface heat transfer coefficient used in GCMs, and [m s-1] is the typical magnitude of the near-surface horizontal winds in GCM simulations (e.g., Lora et al., 2019). In the following sections, we examine how the surface temperature variations reproduced in our GCM simulations at diurnal and seasonal timescales correspond to the theoretical predictions, Eq (6).
2.2 Titan Atmospheric Model
We investigate the diurnal and seasonal variations of surface temperature and PBL height using the Titan Atmospheric Model (TAM), a three-dimensional GCM of Titan that has been extensively validated against observational data (Lora et al., 2015; Lora, 2024). TAM demonstrates particularly strong performance in reproducing lower and middle atmospheric temperatures when compared to observations (Schinder et al., 2011, 2012, 2020; Jennings et al., 2016, 2019). The model optionally incorporates one of two PBL parameterization schemes: a non-local K-profile scheme (Holtslag & Boville, 1993) or the Mellor–Yamad scheme (Mellor & Yamada, 1982). In this study, we employ the latter scheme, consistent with the approach taken by Charnay & Lebonnois (2012).
While TAM has demonstrated robust performance in simulating Titan’s climate system, including the methane cycle (Lora et al., 2015; Lora & Ádámkovics, 2017; Faulk et al., 2020; Battalio et al., 2022; Lora et al., 2022; Lora, 2024), we employ a simplified configuration with dry, flat, and homogeneous surface conditions to facilitate direct comparisons with previous studies and isolate the effects of thermal inertia on surface temperature variations. This simplified approach results in surface temperatures that exceed observed brightness temperatures due to the absence of evaporative cooling (MacKenzie et al., 2019). Consequently, our results should be interpreted as an investigation of temperature and PBL variation mechanisms rather than a representation of conditions on Titan as observed by Cassini.
Our investigation encompasses four simulation cases (Table LABEL:tab:table1). Cases 1, 2, and 3 employ a uniform volumetric heat capacity of [JmK-1], with varying thermal conductivities of , , and [WmK-1], respectively, corresponding to thermal inertia values of , , and [TIU]. Case 4 adopts a spatially varying thermal inertia map based on the “convective lakes” scenario of MacKenzie et al. (2019), which characterizes Titan’s surface using Cassini RADAR observations and classifies it into five terrain types: dunes, hummocky, labyrinth, plains, and lakes, which are assumed to be convecting. In our study, we fix across Cases 1–3 as the volumetric heat capacity values across terrain types in MacKenzie et al. (2019) are relatively similar in magnitude. In contrast, thermal conductivity spans over two orders of magnitude, and our three selected values are chosen to broadly represent this range.
The simulations are run at T21 horizontal resolution (64 × 32 longitude-latitude grid) with 50 atmospheric vertical levels and 15 subsurface vertical levels, using a 600 [s] time step. A shortwave albedo of 0.2 and surface thermal emissivity of 0.95 are adopted. The atmospheric vertical resolution in the lower few kilometers is maintained at several tens of meters to sufficiently resolve boundary layer variations. Subsurface layer depths follow and , where represents the diurnal skin depth for Titan and denotes the subsurface level. This configuration ensures adequate resolution for both diurnal and seasonal subsurface heat conduction by progressively increasing layer depth, with 5 layers within , 10 layers within (the annual skin depth), and 5 layers extending to the deepest region at .
The model is initialized from a cold start with uniform surface and subsurface temperatures of [K], and integrated for 200 Titan years with diurnally averaged output. For diurnal analyses, the simulation is extended by one additional Titan year with output saved every 16 Earth hours (i.e., 1/24 of a Titan day), focusing on a 30 Titan-day period centered around southern midsummer in the final year.
3 Surface Temperature
3.1 Diurnal Variations
Figure 1 presents the diurnal evolution of surface energy balance components and surface temperature near the Huygens landing site during midsummer for the four simulation cases. In Case 4, the plains terrain type ( [TIU]) is assumed at the Huygens landing site (MacKenzie et al., 2019). In Case 1, the net shortwave (SW) flux increases from 0.0 to approximately 1.3 [Wm-2] from nighttime to local noon, while the sensible heat flux peaks near 1.2 [Wm-2] around midday, indicating that a substantial portion of incoming energy is returned to the atmosphere. This behavior reflects the limited capacity of low thermal inertia surfaces to conduct heat into the subsurface, necessitating that excess surface energy be dissipated through sensible heat flux. As thermal inertia increases (Cases 2 and 3), the amplitude of the sensible heat flux correspondingly decreases (1.0 and 0.7 [Wm-2]), as subsurface heat conduction becomes more effective at removing energy from the surface. Meanwhile, the net infrared (IR) radiative flux remains relatively constant at approximately 0.07 [Wm-2] across all cases, consistent with the relatively small changes in surface temperature.
The simulated diurnal surface temperature variations are approximately , , and [K] for Cases 1, 2, and 3, respectively. These values represent the mean and one standard deviation computed over a 30 Titan-day period centered on midsummer in the final simulation year. For a diurnal forcing frequency of [s-1] and net energy flux amplitudes of , , and [Wm-2] for Cases 1, 2, and 3, respectively—derived using the same 30 Titan-day averaging method as for the simulated surface temperature amplitudes—and with an atmospheric damping coefficient of [WmK-1], obtained as the 30-day mean value, the theoretical temperature variations according to Eqs. (6) and (7) are estimated to be , 0.16, and 0.13 [K]. These estimates closely align with the simulated values, reinforcing the validity of the theoretical framework despite its simplification of the net energy flux as a sine-like function. For Case 4, the [Wm-2] yields a theoretical expectation of [K], closely matching the [K] as well.
Despite the fact that the net energy flux—the residual of the surface energy budget—increases with thermal inertia, the amplitude of surface temperature variations decreases. The decreasing amplitude of temperature variation arises because increasing thermal inertia enhances subsurface heat conduction, effectively buffering the surface temperature despite the higher available energy. The values of corresponding to the thermal inertias used in Cases 1, 2, and 3 are approximately 0.37, 1.17, and 3.69 [WmK-1], respectively. These values significantly exceed the atmospheric damping coefficient, [WmK-1], indicating that subsurface heat conduction is the dominant process damping surface temperature variations at diurnal timescales.
3.2 Seasonal Variations
Figure 2 shows the annual zonal-mean surface temperatures as a function of latitude and time for the four simulation cases. As previously reported in MacKenzie et al. (2019)—which explored cases using a uniform thermal inertia of [TIU] (representing porous ice regolith; Tokano, 2005), a globally averaged value of [TIU], and a spatially varying thermal inertia based on the “convective lakes” scenario—seasonal surface temperatures in our Cases 1, 2, and 4 similarly exhibit little to no sensitivity to thermal inertia. In Case 4, the surface is dominated by low thermal inertia terrains: Plains make up approximately 65% of the surface area ( [TIU]), dunes 18% ( [TIU]), and labyrinth terrain 3% ( [TIU]). In contrast, high thermal inertia terrain—hummocky regions with [TIU]—comprise only 14% of the surface (see Figure 2 from MacKenzie et al., 2019). Among the simulation cases, Case 3 alone exhibits a distinct structure compared to the other three cases, indicating that seasonal surface temperature variations begin to show a noticeable sensitivity to thermal inertia only when the thermal inertia is sufficiently high globally.
This behavior can be qualitatively understood using Eqs. (6) and (7). For an annual forcing frequency of [s-1], the corresponding values of for Cases 1, 2, and 3 are approximately 0.014, 0.045, and 0.14 [WmK-1], respectively. Among these, only Case 3 exceeds the atmospheric damping coefficient [WmK-1], obtained as the annual mean, indicating that subsurface heat conduction becomes significant only when thermal inertia is sufficiently high. On seasonal timescales, subsurface conduction is generally less important than atmospheric damping primarily due to Titan’s long orbital period (29.5 Earth years), which yields a small . This stands in contrast to the diurnal regime, wherein subsurface heat conduction can dominate the surface temperature variability at intermediate values of thermal inertia.
Figure 3 presents the seasonal evolution of surface energy balance components and surface temperature near the Huygens landing site for the four simulation cases. Across all cases, the net SW flux varies seasonally from approximately 0.20 to 0.37 [Wm-2], increasing from winter through summer. The sensible heat flux exhibits a broadly similar seasonal pattern, ranging from 0.15 to 0.30 [Wm-2], but with lower thermal inertia cases showing larger fluctuation amplitudes. This enhanced fluctuation arises because, when subsurface heat conduction is inefficient due to low thermal inertia, a stronger sensible heat flux is required to remove the net energy at the surface, as also occurs at diurnal timescales. The net IR flux remains relatively steady at approximately 0.07 [Wm-2] across seasons and simulation cases. The resulting net energy flux is smaller in the lower thermal inertia cases, where a stronger sensible heat flux efficiently removes much of the incoming surface energy.
Surface temperature variations across seasons remain modest—under 1 [K]—and exhibit broadly similar patterns across all simulation cases. However, the amplitude of these variations is more strongly damped for higher thermal inertia. Applying Eqs. (6) and (7) to the simulated seasonal net energy flux amplitudes of , , and [Wm-2] for Cases 1, 2, and 3, respectively, yields theoretical temperature variations of , 0.57, and 0.51 [K]. Here, is derived from means calculated over the final 30 simulation years where 10-Titan-day moving averages are used to remove short-term variations. These estimates are roughly consistent with the simulated surface temperature variations of , , and [K], though some discrepancies remain. These differences likely stem from the assumption of a constant across seasons, whereas in reality the horizontal wind speed exhibits seasonal variability in the simulations. Despite discrepancies between and due to the simplification, Eqs. (6) and (7) effectively capture the dominant role of atmospheric damping in regulating surface temperature variations at seasonal timescales. For instance, omitting the atmospheric damping term leads to theoretical predictions of , 1.22, and 0.70 [K], which significantly overestimate the temperature response compared to the simulations. For Case 4, the [Wm-2] yields a theoretical surface temperature variation of [K], nearly identical to that of Case 2.
4 Planetary Boundary Layer
4.1 Diurnal Variations
Figure 4 presents the diurnal evolution of the PBL height across the four simulation cases. The left panels in each case show the temporal evolution of potential temperature profiles from 9:00 AM to 3:00 PM local solar time, compared with in-situ measurements from the Huygens probe (Fulchignoni et al., 2005). The right panels display the corresponding temperature diffusion tendencies predicted by the model’s PBL parameterization. Consistent with the simulated surface temperature variations of [K] across the cases, temporal changes in the PBL structure are evident in both the potential temperature profiles and diffusion tendencies. The PBL depth increases from approximately 500 [m] in the late morning to about 1,000 [m] by mid-afternoon.
Temperature diffusion tendencies weaken with increasing thermal inertia, consistent with the reduction in sensible heat flux shown in Figure 1. Despite these differences, the simulated PBL heights remain similar across cases, with maximum depths reaching approximately 1,000 [m] regardless of thermal inertia. However, notable differences emerge in the upper troposphere—around 100 [mbar], corresponding to an altitude of approximately 50 [km]—where the simulations diverge. Specifically, Case 1, representing the low thermal inertia scenario, exhibits a higher tropopause temperature of about 71 [K], compared to 70 [K] in Case 3 (high thermal inertia), with Case 2 falling in between (not shown).
This difference reflects the cumulative effect of sensible heat flux over time: low thermal inertia surfaces promote greater upward energy transfer. As a result, the thermal influence extends throughout the troposphere, raising the equilibrium temperature near the tropopause. In contrast, high thermal inertia reduces sensible heat flux, limiting the energy available to warm the free atmosphere. In summary, while thermal structure differences among cases appear minimal at the diurnal timescale, they become more pronounced over longer periods due to the cumulative adjustment of the atmosphere to surface energy exchange. Relatedly, the influence of thermal inertia differences on seasonal wind structures is discussed in the following subsection.
4.2 Seasonal Variations
Figure 5 presents the seasonal evolution of the PBL height across the four simulation cases. The left panels in each case display global zonal means, while the right panels show seasonal variations near the Huygens landing site, alongside the temperature diffusion tendency. The PBL height is diagnosed as the altitude of the model layer at which the vertical diffusion coefficient first drops below a threshold of 0.01 [m2s-1], a criterion commonly employed within the Mellor-Yamada scheme to define the extent of the near-surface adiabatic layer. These diagnosed PBL heights reflect variations in surface heating and the resulting convective mixing, as corroborated by the simulated temperature diffusion tendencies in our simulations.
The global seasonal structure of the PBL closely mirrors the surface temperature patterns shown in Figure 1, indicating that PBL depth variations are primarily governed by the seasonal migration of regions with peak surface heating. The hemispheric asymmetry in the PBL structure is evident, as it is in the surface temperature, due to orbital eccentricity. In Case 1, the diagnosed PBL height at the Huygens landing site varies from approximately 300 [m] during southern winter to about 1,000–1,500 [m] between the southern vernal () and autumnal () equinoxes.
Unlike at diurnal timescales, diffusion tendencies at seasonal timescales exhibit clear sensitivity to thermal inertia: maximum PBL heights reach approximately 2,000 [m] and 3,000 [m] in Cases 2 and 3, respectively. The elevated PBL depth for higher thermal inertia corresponds to a reduced contrast between the minimum and maximum surface temperature, as shown in Figure 3.
Figure 6 presents global seasonal mean maps of the diagnosed PBL depth for four distinct seasons, based on Case 4. The maximum depth is around 1,800 [m], and structural differences across longitudes are apparent due to spatial variations in thermal inertia. For instance, an elevated PBL depth is observed in the Xanadu region (approximately 250°E near the equator) across the seasons, where the assumed thermal inertia corresponds to hummocky terrain (1,962 [TIU]), compared to dune fields (236 [TIU]) and plains (601 [TIU]) at similar latitudes. The imprint of this enhancement is especially pronounced during the southern summer and winter solstices. Additionally, elevated PBL depths are evident in polar regions—near 90°S, 180°E during southern summer and near 70°N, 40°E during southern winter—attributable to the presence of convective lakes with extremely high thermal inertia (18,090 [TIU]). These results suggest that locally varying thermal inertia strongly influences the seasonal evolution of the PBL depth.
Figure 7 shows the zonal and meridional wind profiles at the model grid point closest to the Huygens landing site for the four simulation cases, compared with the Huygens probe observations (Bird et al., 2005; Karkoschka, 2016). The seasonal mean is computed as a 30 Titan-day average centered around midsummer. During the period, the simulated wind profiles exhibit minimal diurnal variation (not shown). Across all cases, the simulations show good agreement with the observed wind structure, capturing key features of both the zonal and meridional components. The simulated near-surface wind magnitudes are on the order of – [ms-1], consistent with results from previous Titan model intercomparison studies (Lora et al., 2019).
In the case of zonal winds, the upper atmosphere on Titan is dominated by eastward (superrotating) flow, driven by the upward and equatorward transport of angular momentum by the mean flow and eddies, respectively (Flasar & Achterberg, 2009; Lewis et al., 2023; Lombardo & Lora, 2023). Near the surface, frictional drag keeps the flow closer to Titan’s solid body rotation. The net gain of angular momentum from the surface at low latitudes necessitates mean westward (retrograde) winds near the surface, a feature consistently reproduced across all four simulations in our study and in prior Titan GCM studies (e.g., Lora et al., 2019). Although differences among simulation cases are minimal near the surface, substantial variations emerge in the stratosphere. In Cases 1, 2, and 4, maximum wind speeds reach approximately 150 [ms-1], whereas in Case 3 they peak at only 60 [ms-1] (not shown). This suggests that differences in global surface temperature structure—driven by varying thermal inertia—can influence wind patterns even at high altitudes. While this implies that surface-induced thermal structures may propagate upward and modulate stratospheric circulation, a detailed analysis of the vertical momentum transfer and dynamical response is beyond the scope of this study.
For meridional winds, the flow near the surface is predominantly southward, transitioning to northward at altitudes around 2 [km]. This vertical wind reversal suggests the presence of a cross-equatorial Hadley circulation, characterized by rising motion in the southern summer hemisphere and descending motion in the northern winter hemisphere. As a result, there is northward return flow aloft and a compensating southward return flow near the surface at the latitude of the Huygens landing site (Lora et al., 2025). Wind magnitudes—both zonal and meridional—are generally smaller in the higher thermal inertia cases, although the differences are minimal. This likely reflects weaker temperature gradients, as higher thermal inertia surfaces dampen surface temperature contrasts (Figure 2), thereby reducing the thermal forcing that drives the atmospheric circulation.
5 PBL Characteristics at the Dragonfly Landing Site
In this section, we present the same analysis of surface temperature and PBL structure as in the previous sections, but for the planned Dragonfly landing site near Selk Crater (E, N) during southern midsummer () (Barnes et al., 2021; Lorenz et al., 2021). For Case 4, the dunes terrain type, with a thermal inertia of [TIU], is assumed at the surface (MacKenzie et al., 2019).
Figure 8 shows the diurnal energy balance components at the Dragonfly landing site. Maximum insolation is reduced compared to the Huygens landing site, as the Dragonfly site experiences northern winter conditions while the Huygens site corresponds to southern summer. This results in a lower mean surface temperature at the Dragonfly site, reduced by approximately 0.1 [K] relative to the Huygens site. The amplitude of diurnal surface temperature variations remains comparable to that at the Huygens site. For Case 4, the simulated diurnal amplitude is [K].
Figure 9 shows the diurnal PBL structures at the Dragonfly landing site. Due to the reduced surface energy input and lower temperatures during this winter season—compared to the midsummer conditions at the Huygens landing site—the diagnosed PBL depths are shallower, reaching a maximum of approximately 500 [m], in contrast to 1,000 [m] at the Huygens site. Nevertheless, the overall temperature structures remain similar to those observed at the Huygens landing site.
Figure 10 shows the seasonal evolution of PBL depth at the Dragonfly landing site. As with the Huygens landing site, higher thermal inertia surfaces lead to elevated PBL depths. However, a key difference is observed in the seasonal symmetry. At the Dragonfly site, the PBL depth exhibits a more symmetric seasonal pattern: maxima occur near the equinoxes ( and ), and minima near the solstices ( and ). In contrast, the Huygens site displays an asymmetric pattern: a pronounced minimum at the southern winter solstice (), with a higher but still suppressed depth at the southern summer solstice (), relative to the equinoctial maxima. This asymmetry is attributed to Saturn’s orbital eccentricity, as illustrated in Figure 2 and the left panels of Figure 5.
Finally, Figure 11 shows the seasonal wind structures at the Dragonfly landing site. Overall, the wind structures remain broadly similar between the Huygens and Dragonfly sites: westward zonal winds dominate the lowest altitudes, with southward flow near the surface and northward flow at higher altitudes. This similarity is expected, as both sites are located near Titan’s equator and are influenced by the large-scale atmospheric circulation. However, subtle differences are apparent. For instance, near-surface zonal winds at the Dragonfly site are more inclined toward negative values compared to those at the Huygens site, consistent with previous results (Lora et al., 2019). The stronger westward near-surface winds at the Dragonfly site, compared to the Huygens site, may be related to its hemispheric position relative to Titan’s seasonally shifting Hadley circulation.
6 Discussion and Conclusions
This study presents a comprehensive reassessment of surface temperature variations on Titan across diurnal and seasonal timescales, and how these variations influence the evolution of the PBL. By combining theoretical analysis with dry GCM simulations, we identify distinct physical mechanisms that govern surface temperature variations at different timescales: subsurface heat conduction dominates the damping of diurnal variations, while atmospheric damping plays a more prominent role at seasonal timescales.
At diurnal timescales, surface temperature amplitudes are highly sensitive to thermal inertia. Low thermal inertia surfaces respond inefficiently to subsurface heat conduction, leading to larger temperature variations and stronger sensible heat fluxes. In contrast, high thermal inertia surfaces store and conduct heat more effectively, resulting in reduced temperature variation amplitude. Theoretical predictions developed in this work accurately reproduce the simulated diurnal temperature variations across different cases, confirming that subsurface heat conduction is the primary damping mechanism at short timescales.
At seasonal timescales, surface temperature variations show much weaker sensitivity to thermal inertia. Only when thermal inertia is uniformly elevated (e.g., Case 3) does seasonal temperature amplitude decline significantly. This is because the product , which characterizes the impact of subsurface heat conduction, is small compared to the atmospheric damping coefficient except under globally high thermal inertia. Consequently, seasonal surface temperature variations are mainly regulated by direct surface–atmosphere energy exchange through sensible heat flux, rather than by subsurface thermal diffusion.
PBL behavior differs markedly between diurnal and seasonal timescales. At diurnal timescales, the PBL exhibits a regular cycle of growth and decay—rising from approximately 500 [m] in the morning to nearly 1,000 [m] in the afternoon—driven by daily variations in surface heating and cooling. At seasonal timescales, the PBL responds to broader shifts in insolation. As thermal inertia increases, the contrast between minimum and maximum surface temperature diminishes, leading the seasonal PBL to extend to a maximum depth of 2,000–3,000 [m], consistent with previous dry GCM simulations (Charnay & Lebonnois, 2012), dune spacing analysis (Lorenz et al., 2010), and the dry adiabatic lapse rate extending to 3.5 [km] observed by Voyager IRIS (Lindal et al., 1983). These results highlight that PBL depth is governed not only by transient energy inputs but also by the long-term thermal state of the surface and overlying atmosphere.
Simulated winds on Titan exhibit minimal diurnal variability, which is probably linked with the small diurnal temperature swings ( [K]) compared to Earth. However, the GCM accurately captures seasonal wind structures, including the key zonal and meridional features observed by the Huygens probe. The simulations support the interpretation by Tokano (2007) that Titan’s wind profiles are shaped primarily by global-scale circulation patterns, not Ekman dynamics as on Earth. A notable seasonal transition in the meridional wind is observed in the simulations: during southern summer, the flow is southward near the surface and northward aloft (Figure 7), while during southern winter, this pattern is reversed, with northward flow near the surface and southward aloft (not shown). This transition is characteristic of Titan’s cross-equatorial Hadley circulation.
At the Dragonfly landing site, simulations during northern midwinter indicate a shallower diurnal PBL (reaching 500 [m]) due to reduced insolation compared to the Huygens site’s midsummer conditions. Nonetheless, temperature structures remain similar, and large-scale wind patterns are preserved, reflecting the shared equatorial latitude and the dominance of global atmospheric dynamics. These results suggest that Dragonfly will likely experience steady boundary layer conditions and modest winds, comparable to those encountered by Huygens.
While our findings offer valuable insights, a key limitation of this study is the exclusion of methane moisture processes, which are likely to play a significant role in Titan’s climate system. Observations from the Huygens probe indicated that the surface was damp at the time of landing (Zarnecki et al., 2005; Lorenz et al., 2006), and our dry simulations overestimate surface temperatures by approximately 2 [K]. Furthermore, in several simulation cases, seasonal maxima in surface temperature and PBL depth occur at high latitudes, contrary to observational evidence indicating that Titan’s warmest regions are consistently located within 30° of the equator, as shown by Voyager IRIS data (Flasar et al., 1981; Flasar, 1998) and 13 years of Cassini CIRS observations (Jennings et al., 2016, 2019). Previous studies have demonstrated that the inclusion of moist processes substantially reduces the pole-to-equator surface temperature contrast, bringing simulated climate patterns into closer agreement with observations (Lora et al., 2019).
This discrepancy highlights the importance of including latent heat processes from methane evaporation and condensation. Prior studies have shown these to play a critical role in regulating Titan’s surface energy balance and atmospheric structure (Griffith et al., 2008; Mitchell et al., 2009; Lora & Ádámkovics, 2017). Incorporating methane-driven moist processes into the surface energy balance term in Eq. (1) will be essential for future modeling efforts. Moreover, Huygens measurements of nearly uniform methane mole fractions in the lowest 4–5 [km] (Niemann et al., 2005, 2010; Gautier et al., 2024) suggest vigorous vertical mixing, possibly linked to the same large-scale circulation that shapes meridional wind reversals. All of this further motivates the inclusion of humidity and moist effects in future Titan PBL simulations.
In summary, this study establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability across diurnal and seasonal timescales. By integrating theoretical energy balance analysis with dry GCM simulations, we demonstrate the important roles of subsurface conduction and atmospheric damping in modulating surface conditions. While future work must incorporate methane-related processes to capture Titan’s full climate complexity, our results provide critical constraints for interpreting Huygens-era observations and offer quantitative projections for environmental conditions at the Dragonfly landing site, supporting mission planning and future data analysis.
Appendix A Theoretical Framework for Surface Energy Balance and Temperature Variability
We seek a solution to the heat diffusion equation:
subject to the boundary conditions:
and
We consider a solution of the form:
where represents the complex amplitude of the temperature variation. The partial derivative of with respect to time is:
Substituting this into (A1) gives:
Assuming is constant (homogeneous subsurface), the equation simplifies to:
Rearranging it, we find:
where . The general solution to this second-order differential equation is . Since the boundary condition (A3) requires the solution to remain finite as , the coefficient must be zero, thus . Using , we find:
At , this becomes:
Applying the boundary condition (A2):
Rearranging gives:
Noting that , the subsurface temperature at any time and depth can be expressed as:
The first and second terms on the right-hand side collectively represent the mean temperature, while the third term denotes the oscillatory component. The factor in the third term describes the exponential decay of the thermal wave amplitude with depth, whereas represents the progressive phase lag between the surface forcing and the thermal response at depth . These terms capture the principal characteristics of heat diffusion in the subsurface medium. The temperature at becomes:
Without the atmospheric coupling coefficient , Eq. (A14) simplifies to the classical solution for surface temperature variations under idealized single-harmonic periodic forcing (Stull, 1988; Garratt, 1992):
This solution reveals a phase lag of between the forcing and surface temperature variations when subsurface heat conduction dominates the surface temperature response. Given that the atmosphere and surface are thermally coupled through radiative and convective processes, this phase lag must also manifest in the atmospheric response. Consequently, we can express the atmospheric coupling coefficient as . Substituting this expression into Eq. (A14) yields:
The amplitude of surface temperature oscillations is given by:
where [Wm-2] denotes the amplitude of fluctuations in . This expression is notably independent of the specific value of the deep boundary condition , ensuring its general applicability.

### v2

@@
 Understanding Titan’s planetary boundary layer (PBL)—the lowest region of the atmosphere influenced by surface conditions—remains challenging due to Titan’s thick atmosphere and limited observations.
-Previous modeling studies have also produced inconsistent estimates of surface temperature, a critical determinant of PBL behavior, often without clear explanations grounded in surface energy balance.
-In this study, we develop a theoretical framework and apply a three-dimensional dry general circulation model (GCM) to investigate how surface thermal inertia influences surface energy balance and temperature variability across diurnal and seasonal timescales.
-At diurnal timescales, lower thermal inertia surfaces exhibit larger temperature swings and enhanced sensible heat fluxes due to inefficient subsurface heat conduction.
+Previous modeling studies have also produced inconsistent estimates of surface temperature variability, a critical determinant of PBL behavior, often without clear explanations grounded in surface energy balance.
+Here, we develop a theoretical framework and apply a three-dimensional dry general circulation model (GCM) to investigate how surface thermal inertia influences surface energy balance and temperature variability across diurnal and seasonal timescales.
+At diurnal timescales, lower thermal inertia surfaces experience larger temperature fluctuations and enhanced daytime sensible heat fluxes due to less efficient subsurface heat conduction.
 In contrast, at seasonal timescales, surface temperature variations show weak sensitivity to thermal inertia, as atmospheric damping tends to dominate over subsurface conduction.
-The PBL depth ranges from a few hundred meters to 1,000 [m] on diurnal timescales, while seasonal maxima reach 2,000–3,000 [m], supporting the interpretation from a previous study that the Huygens probe captured the two PBL structures.
+The PBL depth ranges from a few hundred meters to 1,000 m on diurnal timescales, while seasonal maxima reach 2,000 m, supporting the interpretation from a previous study that the Huygens probe captured the two PBL structures.
 Simulated seasonal winds at the Huygens landing site successfully reproduce key observed features, including near-surface retrograde winds and meridional wind reversals within the lowest few kilometers, consistent with Titan’s cross-equatorial Hadley circulation.
-Simulations at the planned Dragonfly landing site predict shallower thermal PBLs with broadly similar wind patterns.
+Simulations for the planned Dragonfly landing site predict shallower thermal PBLs with smaller fluctuation amplitudes, while maintaining similar wind patterns.
 This work establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability, and offers a unified explanation of Titan’s PBL behavior that provides improved guidance for future missions.
@@
 These transitions between temperature extrema, whether diurnal or seasonal, generate distinct atmospheric stability regimes that influence critical weather and climate processes, including cloud formation, precipitation, local breezes, and thunderstorms (Stull, 1988; Garratt, 1992; Lee, 2016).
-The phenomenon of the PBL extends beyond Earth to other celestial bodies possessing both a solid surface and an atmosphere (Petrosyan et al., 2011; Lebonnois et al., 2013; Flasar et al., 2014).
+The phenomenon of the PBL extends beyond Earth to other celestial bodies possessing both a solid surface and an atmosphere (Petrosyan et al., 2011; Flasar et al., 2014; Lebonnois et al., 2018).
 As such, studying the PBL of other planetary bodies provides valuable insights into general PBL processes, making it crucial for comparative planetology.
@@
 These striking geographical contrasts make understanding Titan’s PBL essential for explaining the formation and evolution of such diverse landscapes.
-Due to Titan’s vast distance from the Sun and its thick, opaque atmosphere, average surface insolation is estimated at approximately 0.1 % of Earth’s, around 1.0 [Wm-2] (McKay et al., 1989, 1991; Tomasko et al., 2008).
+Due to Titan’s vast distance from the Sun and its thick, opaque atmosphere, average surface insolation is estimated at approximately 0.1 % of Earth’s, around 1.0 Wm-2 (McKay et al., 1989, 1991; Tomasko et al., 2008).
 However, given Titan’s extended day (16 Earth days) and year (29.5 Earth years), the accumulated surface energy over these cycles may suffice to generate notable surface temperature variations and PBL responses.
@@
 Understanding how surface temperature responds to solar forcing is essential for interpreting variations in the PBL.
-Within the context of modeling surface temperature variations on terrestrial planets, thermal inertia, defined as [JmsK-1] (hereafter we will use thermal inertia units; TIU), serves as a crucial parameter for quantifying a surface’s thermal response to energy input (Price, 1977; Putzig et al., 2005; MacKenzie et al., 2019).
-Here, represents the density [kgm-3], denotes the specific heat capacity [JkgK-1], and is the thermal conductivity [WmK-1] of the surface.
+Within the context of modeling surface temperature variations on terrestrial planets, thermal inertia, defined as (JmsK-1; hereafter we will use thermal inertia units; TIU), serves as a crucial parameter for quantifying a surface’s thermal response to energy input (Price, 1977; Putzig et al., 2005; MacKenzie et al., 2019).
+Here, represents the density (kgm-3), denotes the specific heat capacity (JkgK-1), and is the thermal conductivity (WmK-1) of the surface.
 Materials with lower thermal inertia tend to exhibit larger temperature variations, reflecting their diminished capacity to store and conduct heat (Stull, 1988; Garratt, 1992).
-Using a dry general circulation model (GCM) of Titan’s atmosphere coupled with a simplified subsurface scheme, Tokano (2005) reported diurnal and seasonal surface temperature variations of a few kelvins near the equator—close to the latitude of the Huygens landing at 10.3°S—assuming a uniform thermal inertia of [TIU], characteristic of porous ice regolith.
+Using a dry general circulation model (GCM) of Titan’s atmosphere coupled with a simplified subsurface scheme, Tokano (2005) reported diurnal and seasonal surface temperature variations of a few kelvins near the equator—close to the latitude of the Huygens landing at 10.3°S—assuming a uniform thermal inertia of TIU, characteristic of porous ice regolith.
 Williams et al.
 (2012) employed a one-dimensional energy-balance model, considering the Huygens landing site with a fixed near-surface atmospheric temperature, to investigate surface temperature variations.
-Their results indicated smaller diurnal and seasonal surface temperature variations, with amplitudes less than 0.1 [K] under dry conditions and a thermal inertia of [TIU], contradicting the findings of Tokano (2005).
-Subsequently, Charnay & Lebonnois (2012) reported diurnal fluctuations around 0.3 [K] in surface temperatures at the Huygens landing site using a dry GCM with [TIU] (Lebonnois et al., 2012).
-They proposed that the Huygens measurement of a near-adiabatic layer reaching 300 [m] (Tokano et al., 2006) reflected ongoing convection during the probe’s descent (at approximately 10 am local time), while changes in the lapse rate at 800 [m] and 2,000 [m] corresponded to residual signatures of maximum diurnal PBL extension and seasonal Hadley circulation effects, respectively.
+Their results indicated smaller diurnal and seasonal surface temperature variations, with amplitudes less than 0.1 K under dry conditions and a thermal inertia of TIU, contradicting the findings of Tokano (2005).
+Subsequently, Charnay & Lebonnois (2012) reported diurnal fluctuations around 0.3 K in surface temperatures at the Huygens landing site using a dry GCM with TIU (Lebonnois et al., 2012).
+They proposed that the Huygens measurement of a near-adiabatic layer reaching 300 m (Tokano et al., 2006) reflected ongoing convection during the probe’s descent (at approximately 10 am local time), while changes in the lapse rate at 800 m and 2,000 m corresponded to residual signatures of maximum diurnal PBL extension and seasonal Hadley circulation effects, respectively.
 More recently, MacKenzie et al.
@@
 Section 5 evaluates expected conditions at the Dragonfly landing site.
-Section 6 discusses the implications of the findings and and concludes by highlighting the distinct roles of atmospheric damping and subsurface heat conduction in regulating surface temperature and PBL structures—dominant at seasonal and diurnal timescales, respectively.
+Section 6 discusses the implications of the findings and concludes by highlighting the distinct roles of atmospheric damping and subsurface heat conduction in regulating surface temperature and PBL structures—dominant at seasonal and diurnal timescales, respectively.
 Overall, this study provides a critical reassessment of the atmospheric profile measured by Huygens and offers new insights into Titan’s lower atmosphere, informing expectations for the upcoming Dragonfly mission.
@@
 Energy balance at the surface for a dry case (i.e., without latent heat flux) is given by:
-where and [Wm-2] represent the net radiative flux into and the sensible heat flux out of the surface, respectively.
-The residual, [Wm-2], is the net energy flux at the surface available for subsurface heat conduction.
+where and (Wm-2) represent the net radiative flux into and the sensible heat flux out of the surface, respectively.
+The residual, (Wm-2), is the net energy flux at the surface available for subsurface heat conduction.
 The evolution of surface temperature follows the subsurface heat diffusion equation using as the surface boundary condition:
-where [JmK-1] collectively denotes volumetric heat capacity.
+where (JmK-1) collectively denotes volumetric heat capacity.
 Here, homogeneous subsurface properties are considered for simplicity and to compare with previous results.
 The net energy flux at the surface can be simplified as a single harmonic oscillator with linear atmospheric damping (e.g., Dickinson, 1988):
-where [Wm-2] is the amplitude of the periodic flux, and [Wm-2K-1] is the atmospheric damping coefficient.
+where is the amplitude of the periodic flux, and (Wm-2K-1) is the atmospheric damping coefficient.
 Over land surfaces, as assumed in dry GCM configurations, an additional lower boundary condition must be applied to the subsurface; for example:
-where [K] is some mean temperature in the deep subsurface, and [m] is the thermal skin depth corresponding to a forcing frequency of [s-1], whether it be diurnal or annual.
+where is some mean temperature in the deep subsurface, and (m) is the thermal skin depth corresponding to a forcing frequency of (s-1), whether it be diurnal or annual.
 The magnitude of surface temperature fluctuations is expressed as (see Appendix A for the full derivation):
-where [Wm-2] denotes the amplitude of fluctuations in .
-The thermal inertia [TIU], combined with [s-0.5] represents damping of the energy ([Wm-2K-1]) received at the surface, due to subsurface heat conduction.
-Meanwhile, [Wm-2K-1], which is the magnitude of , accounts for damping due to direct energy exchange between the surface and the atmosphere.
-For Titan, the temperature of the overlying atmosphere, [K], can be reasonably assumed to remain constant due to its long radiative timescale (Flasar et al., 1981, 2014).
+where denotes the amplitude of fluctuations in .
+The thermal inertia , combined with represents damping of the energy (Wm-2K-1) received at the surface, due to subsurface heat conduction.
+Meanwhile, , which is the magnitude of , accounts for damping due to direct energy exchange between the surface and the atmosphere.
+For Titan, the temperature of the overlying atmosphere, , can be reasonably assumed to remain constant due to its long radiative timescale (Flasar et al., 1981, 2014).
 Notably, the linear atmospheric damping term in Eq.
 (4) takes the same form as the expression for sensible heat flux:
-where [kgm-3] is the density of Titan’s lower atmosphere as measured by the Huygens probe, is the specific heat capacity of the N2-dominated atmosphere [JkgK-1], [dimensionless] is the surface heat transfer coefficient used in GCMs, and [m s-1] is the typical magnitude of the near-surface horizontal winds in GCM simulations (e.g., Lora et al., 2019).
+where (kgm-3) is the density of Titan’s lower atmosphere as measured by the Huygens probe, is the specific heat capacity of the N2-dominated atmosphere (JkgK-1), is the surface heat transfer coefficient used in GCMs, and (m s-1) is the typical magnitude of the near-surface horizontal winds in GCM simulations (e.g., Lora et al., 2019).
 In the following sections, we examine how the surface temperature variations reproduced in our GCM simulations at diurnal and seasonal timescales correspond to the theoretical predictions, Eq (6).
@@
 Our investigation encompasses four simulation cases (Table LABEL:tab:table1).
-Cases 1, 2, and 3 employ a uniform volumetric heat capacity of [JmK-1], with varying thermal conductivities of , , and [WmK-1], respectively, corresponding to thermal inertia values of , , and [TIU].
+Cases 1, 2, and 3 employ a uniform volumetric heat capacity of JmK-1, with varying thermal conductivities of , , and WmK-1, respectively, corresponding to thermal inertia values of , , and TIU.
 Case 4 adopts a spatially varying thermal inertia map based on the “convective lakes” scenario of MacKenzie et al.
@@
 In contrast, thermal conductivity spans over two orders of magnitude, and our three selected values are chosen to broadly represent this range.
-The simulations are run at T21 horizontal resolution (64 × 32 longitude-latitude grid) with 50 atmospheric vertical levels and 15 subsurface vertical levels, using a 600 [s] time step.
+The simulations are run at T21 horizontal resolution (64 × 32 longitude-latitude grid) with 50 atmospheric vertical levels and 15 subsurface vertical levels, using a 600 s time step.
 A shortwave albedo of 0.2 and surface thermal emissivity of 0.95 are adopted.
+The radiative transfer parameterization in TAM, which uses plane-parallel two stream approximations, is non-gray and accounts for multiple scattering, employing absorption coefficients and haze optical properties from observations (Lora et al., 2015) and producing radiative heating rates that closely match estimates based on Huygens measurements (Tomasko et al., 2008).
+Still, the shortwave heating at high zenith angles (dawn and dusk) should be regarded with caution given the plane parallel assumption, though this issue is reduced in the thick and strongly scattering lower atmosphere.
+Near the equator, where the Huygens and Dragonfly landing sites are located, the horizontal resolution is about 253 km in both latitude and longitude, corresponding to an area of approximately km2 per grid cell.
 The atmospheric vertical resolution in the lower few kilometers is maintained at several tens of meters to sufficiently resolve boundary layer variations.
@@
 This configuration ensures adequate resolution for both diurnal and seasonal subsurface heat conduction by progressively increasing layer depth, with 5 layers within , 10 layers within (the annual skin depth), and 5 layers extending to the deepest region at .
-The model is initialized from a cold start with uniform surface and subsurface temperatures of [K], and integrated for 200 Titan years with diurnally averaged output.
+The model is initialized from a cold start with uniform surface and subsurface temperatures of K, and integrated for 200 Titan years with diurnally averaged output.
 For diurnal analyses, the simulation is extended by one additional Titan year with output saved every 16 Earth hours (i.e., 1/24 of a Titan day), focusing on a 30 Titan-day period centered around southern midsummer in the final year.
@@
 Figure 1 presents the diurnal evolution of surface energy balance components and surface temperature near the Huygens landing site during midsummer for the four simulation cases.
-In Case 4, the plains terrain type ( [TIU]) is assumed at the Huygens landing site (MacKenzie et al., 2019).
-In Case 1, the net shortwave (SW) flux increases from 0.0 to approximately 1.3 [Wm-2] from nighttime to local noon, while the sensible heat flux peaks near 1.2 [Wm-2] around midday, indicating that a substantial portion of incoming energy is returned to the atmosphere.
+In Case 4, the plains terrain type ( TIU) is assumed at the Huygens landing site (MacKenzie et al., 2019).
+In Case 1, the net shortwave (SW) flux increases from 0.0 to approximately 1.3 Wm-2 from nighttime to local noon, while the sensible heat flux peaks near 1.2 Wm-2 around midday, indicating that a substantial portion of incoming energy is returned to the atmosphere.
 This behavior reflects the limited capacity of low thermal inertia surfaces to conduct heat into the subsurface, necessitating that excess surface energy be dissipated through sensible heat flux.
-As thermal inertia increases (Cases 2 and 3), the amplitude of the sensible heat flux correspondingly decreases (1.0 and 0.7 [Wm-2]), as subsurface heat conduction becomes more effective at removing energy from the surface.
-Meanwhile, the net infrared (IR) radiative flux remains relatively constant at approximately 0.07 [Wm-2] across all cases, consistent with the relatively small changes in surface temperature.
-The simulated diurnal surface temperature variations are approximately , , and [K] for Cases 1, 2, and 3, respectively.
-These values represent the mean and one standard deviation computed over a 30 Titan-day period centered on midsummer in the final simulation year.
-For a diurnal forcing frequency of [s-1] and net energy flux amplitudes of , , and [Wm-2] for Cases 1, 2, and 3, respectively—derived using the same 30 Titan-day averaging method as for the simulated surface temperature amplitudes—and with an atmospheric damping coefficient of [WmK-1], obtained as the 30-day mean value, the theoretical temperature variations according to Eqs.
-(6) and (7) are estimated to be , 0.16, and 0.13 [K].
+During the evening and nighttime, the sensible heat flux turns slightly negative, indicating rapid surface cooling and the formation of a near-surface temperature inversion.
+The inversion is evident in the potential-temperature structure and is discussed in Section 4.1.
+As thermal inertia increases (Cases 2 and 3), the daytime sensible heat flux amplitude decreases to 1.0 and 0.8 Wm-2, as subsurface heat conduction more effectively removes energy from the surface.
+During the evening and nighttime, the sensible heat flux shifts toward positive values, reflecting slower surface cooling.
+Meanwhile, the net infrared (IR) radiative flux remains relatively constant at approximately 0.07 Wm-2 across all cases, consistent with the relatively small changes in surface temperature.
+Case 4 shows similar behavior to Case 2 due to their comparable thermal inertias.
+The simulated diurnal surface temperature variations are approximately , , and K for Cases 1, 2, and 3, respectively.
+These values represent the mean and one standard deviation computed over a 30 Titan-day period centered on midsummer (–) in the final simulation year.
+For a diurnal forcing frequency of s-1 and net energy flux amplitudes of , , and Wm-2 for Cases 1, 2, and 3, respectively—derived using the same 30 Titan-day averaging method as for the simulated surface temperature amplitudes—and with an atmospheric damping coefficient of [WmK-1], obtained as the 30-day mean value, the theoretical temperature variations according to Eqs.
+(6) and (7) are estimated to be , 0.16, and 0.13 K.
 These estimates closely align with the simulated values, reinforcing the validity of the theoretical framework despite its simplification of the net energy flux as a sine-like function.
-For Case 4, the [Wm-2] yields a theoretical expectation of [K], closely matching the [K] as well.
+For Case 4, the Wm-2 yields a theoretical expectation of K, closely matching the K as well.
 Despite the fact that the net energy flux—the residual of the surface energy budget—increases with thermal inertia, the amplitude of surface temperature variations decreases.
 The decreasing amplitude of temperature variation arises because increasing thermal inertia enhances subsurface heat conduction, effectively buffering the surface temperature despite the higher available energy.
-The values of corresponding to the thermal inertias used in Cases 1, 2, and 3 are approximately 0.37, 1.17, and 3.69 [WmK-1], respectively.
-These values significantly exceed the atmospheric damping coefficient, [WmK-1], indicating that subsurface heat conduction is the dominant process damping surface temperature variations at diurnal timescales.
+The values of corresponding to the thermal inertias used in Cases 1, 2, 3 and 4 are approximately 0.37, 1.17, 3.69 and 1.28 WmK-1, respectively.
+These values significantly exceed the atmospheric damping coefficient, WmK-1, indicating that subsurface heat conduction is the dominant process in regulating surface temperature variations at diurnal timescales.
 3.2 Seasonal Variations
@@
 As previously reported in MacKenzie et al.
-(2019)—which explored cases using a uniform thermal inertia of [TIU] (representing porous ice regolith; Tokano, 2005), a globally averaged value of [TIU], and a spatially varying thermal inertia based on the “convective lakes” scenario—seasonal surface temperatures in our Cases 1, 2, and 4 similarly exhibit little to no sensitivity to thermal inertia.
-In Case 4, the surface is dominated by low thermal inertia terrains: Plains make up approximately 65% of the surface area ( [TIU]), dunes 18% ( [TIU]), and labyrinth terrain 3% ( [TIU]).
-In contrast, high thermal inertia terrain—hummocky regions with [TIU]—comprise only 14% of the surface (see Figure 2 from MacKenzie et al., 2019).
-Among the simulation cases, Case 3 alone exhibits a distinct structure compared to the other three cases, indicating that seasonal surface temperature variations begin to show a noticeable sensitivity to thermal inertia only when the thermal inertia is sufficiently high globally.
+(2019)—which explored cases using a uniform thermal inertia of TIU (representing porous ice regolith; Tokano, 2005), a globally averaged value of TIU, and a spatially varying thermal inertia based on the “convective lakes” scenario—seasonal surface temperatures in our Cases 1, 2, and 4 similarly exhibit little to no sensitivity to thermal inertia.
+In Case 4, the surface is dominated by low thermal inertia terrains: Plains make up approximately 65% of the surface area ( TIU), dunes 18% ( TIU), and labyrinth terrain 3% ( TIU).
+In contrast, high thermal inertia terrain—hummocky regions with TIU—comprise only 14% of the surface (see Figure 2 from MacKenzie et al., 2019).
+Among our simulation cases, Case 3 alone exhibits a distinct structure compared to the other three, indicating that seasonal surface temperature variations begin to show a noticeable sensitivity to thermal inertia only when the thermal inertia is sufficiently high globally.
 This behavior can be qualitatively understood using Eqs.
 (6) and (7).
-For an annual forcing frequency of [s-1], the corresponding values of for Cases 1, 2, and 3 are approximately 0.014, 0.045, and 0.14 [WmK-1], respectively.
-Among these, only Case 3 exceeds the atmospheric damping coefficient [WmK-1], obtained as the annual mean, indicating that subsurface heat conduction becomes significant only when thermal inertia is sufficiently high.
+For an annual forcing frequency of s-1, the corresponding values of for Cases 1, 2, 3, and 4 are approximately 0.014, 0.045, 0.14, and 0.049 WmK-1, respectively.
+Among these, only Case 3 exceeds the atmospheric damping coefficient WmK-1, obtained as the annual mean, indicating that subsurface heat conduction becomes significant only when thermal inertia is sufficiently high.
 On seasonal timescales, subsurface conduction is generally less important than atmospheric damping primarily due to Titan’s long orbital period (29.5 Earth years), which yields a small .
@@
 Figure 3 presents the seasonal evolution of surface energy balance components and surface temperature near the Huygens landing site for the four simulation cases.
-Across all cases, the net SW flux varies seasonally from approximately 0.20 to 0.37 [Wm-2], increasing from winter through summer.
-The sensible heat flux exhibits a broadly similar seasonal pattern, ranging from 0.15 to 0.30 [Wm-2], but with lower thermal inertia cases showing larger fluctuation amplitudes.
-This enhanced fluctuation arises because, when subsurface heat conduction is inefficient due to low thermal inertia, a stronger sensible heat flux is required to remove the net energy at the surface, as also occurs at diurnal timescales.
-The net IR flux remains relatively steady at approximately 0.07 [Wm-2] across seasons and simulation cases.
+Across all cases, the net SW flux varies seasonally from approximately 0.20 to 0.37 Wm-2, increasing from winter through summer.
+The sensible heat flux exhibits a broadly similar seasonal pattern, ranging from 0.15 to 0.30 Wm-2.
+Lower-thermal-inertia cases exhibit larger fluctuation amplitudes, driven by greater surface temperature variability.
+The net IR flux remains relatively steady at approximately 0.07 Wm-2 across seasons and simulation cases.
 The resulting net energy flux is smaller in the lower thermal inertia cases, where a stronger sensible heat flux efficiently removes much of the incoming surface energy.
-Surface temperature variations across seasons remain modest—under 1 [K]—and exhibit broadly similar patterns across all simulation cases.
+Surface temperature variations across seasons remain modest—under 1 K—and exhibit broadly similar patterns across all simulation cases.
 However, the amplitude of these variations is more strongly damped for higher thermal inertia.
 Applying Eqs.
-(6) and (7) to the simulated seasonal net energy flux amplitudes of , , and [Wm-2] for Cases 1, 2, and 3, respectively, yields theoretical temperature variations of , 0.57, and 0.51 [K].
+(6) and (7) to the simulated seasonal net energy flux amplitudes of , , and Wm-2 for Cases 1, 2, and 3, respectively, yields theoretical temperature variations of , 0.57, and 0.51 K.
 Here, is derived from means calculated over the final 30 simulation years where 10-Titan-day moving averages are used to remove short-term variations.
-These estimates are roughly consistent with the simulated surface temperature variations of , , and [K], though some discrepancies remain.
+These estimates are roughly consistent with the simulated surface temperature variations of , , and K, though some discrepancies remain.
 These differences likely stem from the assumption of a constant across seasons, whereas in reality the horizontal wind speed exhibits seasonal variability in the simulations.
@@
 (6) and (7) effectively capture the dominant role of atmospheric damping in regulating surface temperature variations at seasonal timescales.
-For instance, omitting the atmospheric damping term leads to theoretical predictions of , 1.22, and 0.70 [K], which significantly overestimate the temperature response compared to the simulations.
-For Case 4, the [Wm-2] yields a theoretical surface temperature variation of [K], nearly identical to that of Case 2.
+For instance, omitting the atmospheric damping term leads to theoretical predictions of , 1.22, and 0.70 K, which significantly overestimate the temperature response compared to the simulations.
+For Case 4, the Wm-2 yields a theoretical surface temperature variation of K, nearly identical to that of Case 2.
 4 Planetary Boundary Layer
 4.1 Diurnal Variations
-Figure 4 presents the diurnal evolution of the PBL height across the four simulation cases.
-The left panels in each case show the temporal evolution of potential temperature profiles from 9:00 AM to 3:00 PM local solar time, compared with in-situ measurements from the Huygens probe (Fulchignoni et al., 2005).
-The right panels display the corresponding temperature diffusion tendencies predicted by the model’s PBL parameterization.
-Consistent with the simulated surface temperature variations of [K] across the cases, temporal changes in the PBL structure are evident in both the potential temperature profiles and diffusion tendencies.
-The PBL depth increases from approximately 500 [m] in the late morning to about 1,000 [m] by mid-afternoon.
-Temperature diffusion tendencies weaken with increasing thermal inertia, consistent with the reduction in sensible heat flux shown in Figure 1.
-Despite these differences, the simulated PBL heights remain similar across cases, with maximum depths reaching approximately 1,000 [m] regardless of thermal inertia.
-However, notable differences emerge in the upper troposphere—around 100 [mbar], corresponding to an altitude of approximately 50 [km]—where the simulations diverge.
-Specifically, Case 1, representing the low thermal inertia scenario, exhibits a higher tropopause temperature of about 71 [K], compared to 70 [K] in Case 3 (high thermal inertia), with Case 2 falling in between (not shown).
-This difference reflects the cumulative effect of sensible heat flux over time: low thermal inertia surfaces promote greater upward energy transfer.
-As a result, the thermal influence extends throughout the troposphere, raising the equilibrium temperature near the tropopause.
-In contrast, high thermal inertia reduces sensible heat flux, limiting the energy available to warm the free atmosphere.
-In summary, while thermal structure differences among cases appear minimal at the diurnal timescale, they become more pronounced over longer periods due to the cumulative adjustment of the atmosphere to surface energy exchange.
-Relatedly, the influence of thermal inertia differences on seasonal wind structures is discussed in the following subsection.
+Figure 4 shows the diurnal evolution of PBL height across the four simulation cases near the Huygens landing site.
+In each case, the left panels present snapshots of potential temperature profiles at 9:00 AM, 12:00 PM, and 3:00 PM local solar time (blue), as well as 12:00 AM (purple), compared with in-situ Huygens probe measurements (Fulchignoni et al., 2005) taken around 10:00 AM (black).
+Surface heating and the deepening of the adiabatic layer due to dry convection during the day are evident in all cases.
+At 12:00 AM, a stable near-surface inversion forms in Case 1, consistent with the negative sensible-heat flux in Figure 1.
+As thermal inertia increases across cases, the 12:00 AM profiles become progressively less stable, consistent with reduced nocturnal cooling and a warmer surface.
+Notably, in Case 3 the 12:00 AM profile is similar to those at other local times, indicating that a near-surface adiabatic layer persists throughout the diurnal cycle under high thermal inertia.
+The middle panels show the corresponding temperature diffusion tendencies predicted by the model’s PBL parameterization, with arrows marking the profiles at 9:00 AM, 12:00 PM, and 3:00 PM corresponding to the inflection points in the left panels.
+The diffusion tendencies are generally positive, reflecting upward heat transfer by convection, but contain interspersed negative layers that create a noisy appearance.
+The evolution of these tendencies closely tracks the changes in the potential temperature profiles.
+The right panels display the mean composite evolution of PBL depth, averaged over 30 Titan days during midsummer (–).
+We diagnose PBL height as the height above ground where the vertical eddy diffusivity first falls below .
+This diffusivity-threshold criterion is the default in TAM when the Mellor–Yamada scheme is used and it closely delineates the top of the near-surface adiabatic layer.
+The time series is smoothed with a Savitzky–Golay filter (window = 5 bins, polynomial order = 2) to suppress bin-to-bin noise while preserving diurnal phase and amplitude.
+Results are insensitive to reasonable choices of window (3–7 bins).
+These diagnosed PBL heights reflect variations in surface heating and the resulting convective mixing, as corroborated by the inflection point of the potential temperature profile and the temperature diffusion tendencies in the left and middle panels.
+Although the individual PBL-depth time series show variability with occasional outliers (e.g., Case 2), these are attributable to natural variability; averaging effectively suppresses them.
+For the lowest thermal inertia surface (Case 1), the PBL grows from m to m by mid-afternoon, where uncertainty ranges correspond to one standard deviation from the mean.
+With increasing thermal inertia, baseline nighttime PBL depths rise to m and m, and maximum daytime depths reach m and m for Case 2 and Case 3, respectively.
+The significant increase in nighttime PBL depths for higher thermal inertia cases reflects reduced nighttime surface cooling, which allows the PBL to remain deeper (Figure 1).
+For Case 4, the PBL grows from m to m by mid-afternoon, closely matching Case 2, which employs a comparable thermal inertia.
+Daytime maximum PBL depths are deeper for higher-thermal-inertia surfaces, despite the smaller daytime sensible heat fluxes over these surfaces.
+This likely reflects that daytime depth is controlled by the integrated heat accumulation over the day rather than by instantaneous fluxes.
+Indeed, the PBL fluctuation amplitude—reflecting the magnitude of PBL growth—is greatest for the lowest thermal inertia case, consistent with its larger sensible heat flux variability.
+Differences in sensible heat flux lead to noticeable variations near the tropopause (100 mbar, 50 km altitude): Case 1 shows a higher temperature (71 K) than Case 3 (70 K), with Case 2 and 4 intermediate (not shown).
+This reflects the cumulative effect of sensible heat flux over time: low thermal inertia surfaces transfer more energy upward, warming the free atmosphere and raising tropopause temperature, whereas high thermal inertia surfaces limit this warming.
 4.2 Seasonal Variations
 Figure 5 presents the seasonal evolution of the PBL height across the four simulation cases.
-The left panels in each case display global zonal means, while the right panels show seasonal variations near the Huygens landing site, alongside the temperature diffusion tendency.
-The PBL height is diagnosed as the altitude of the model layer at which the vertical diffusion coefficient first drops below a threshold of 0.01 [m2s-1], a criterion commonly employed within the Mellor-Yamada scheme to define the extent of the near-surface adiabatic layer.
-These diagnosed PBL heights reflect variations in surface heating and the resulting convective mixing, as corroborated by the simulated temperature diffusion tendencies in our simulations.
+The left panels in each case display global zonal means compiled over 10 Titan years, while the right panels show the 10-Titan-year mean composite PBL depth near the Huygens landing site, where Savitzky–Golay smoothing (30 Titan-day window, third order) is applied to emphasize the seasonal amplitude.
 The global seasonal structure of the PBL closely mirrors the surface temperature patterns shown in Figure 2, indicating that PBL depth variations are primarily governed by the seasonal migration of regions with peak surface heating.
 The hemispheric asymmetry in the PBL structure is evident, as it is in the surface temperature, due to orbital eccentricity.
-In Case 1, the diagnosed PBL height at the Huygens landing site varies from approximately 300 [m] during southern winter to about 1,000–1,500 [m] between the southern vernal () and autumnal () equinoxes.
-Unlike at diurnal timescales, diffusion tendencies at seasonal timescales exhibit clear sensitivity to thermal inertia: maximum PBL heights reach approximately 2,000 [m] and 3,000 [m] in Cases 2 and 3, respectively.
-The elevated PBL depth for higher thermal inertia corresponds to a reduced contrast between the minimum and maximum surface temperature, as shown in Figure 3.
-Figure 6 presents global seasonal mean maps of the diagnosed PBL depth for four distinct seasons, based on Case 4.
-The maximum depth is around 1,800 [m], and structural differences across longitudes are apparent due to spatial variations in thermal inertia.
-For instance, an elevated PBL depth is observed in the Xanadu region (approximately 250°E near the equator) across the seasons, where the assumed thermal inertia corresponds to hummocky terrain (1,962 [TIU]), compared to dune fields (236 [TIU]) and plains (601 [TIU]) at similar latitudes.
+In Case 1, the diagnosed PBL height at the Huygens landing site ranges from approximately 180 m during southern winter to about 1,100 m between the southern vernal () and autumnal () equinoxes.
+In Case 2, it ranges from 200 m to 1,320 m, while in Case 3, it extends from 290 m to 2,300 m.
+The increase in baseline (minimum) PBL depths for higher thermal inertia cases is modest, whereas the increase in maximum depths is more pronounced.
+In both cases, the elevated PBL heights for higher thermal inertia simulations are consistent with the diurnal-timescale behavior: a reduced contrast between minimum and maximum surface temperatures produces a more steady temperature structure that allows greater energy buildup and formation of a deeper PBL.
+Figure 6 presents global seasonal mean maps of the diagnosed PBL depth for four distinct seasons, using Case 4 to illustrate how local variations in thermal inertia influence PBL depth.
+The maximum depth is around 1,800 m, and structural differences across longitudes are apparent due to spatial variations in thermal inertia.
+For instance, an elevated PBL depth is observed in the Xanadu region (approximately 250°E near the equator) across the seasons, where the assumed thermal inertia corresponds to hummocky terrain (1,962 TIU), compared to dune fields (236 TIU) and plains (601 TIU) at similar latitudes elsewhere.
 The imprint of this enhancement is especially pronounced during the southern summer and winter solstices.
-Additionally, elevated PBL depths are evident in polar regions—near 90°S, 180°E during southern summer and near 70°N, 40°E during southern winter—attributable to the presence of convective lakes with extremely high thermal inertia (18,090 [TIU]).
+Additionally, elevated PBL depths are evident in polar regions—near 90°S, 180°E during southern summer and near 70°N, 40°E during southern winter, attributable to the presence of “convective lakes” with extremely high thermal inertia (18,090 TIU) (MacKenzie et al., 2019).
 These results suggest that locally varying thermal inertia strongly influences the seasonal evolution of the PBL depth.
 Figure 7 shows the zonal and meridional wind profiles at the model grid point closest to the Huygens landing site for the four simulation cases, compared with the Huygens probe observations (Bird et al., 2005; Karkoschka, 2016).
+Positive values in the zonal and meridional directions correspond to eastward and northward motions, respectively.
 The seasonal mean is computed as a 30 Titan-day average centered around midsummer.
 During the period, the simulated wind profiles exhibit minimal diurnal variation (not shown).
-Across all cases, the simulations show good agreement with the observed wind structure, capturing key features of both the zonal and meridional components.
-The simulated near-surface wind magnitudes are on the order of – [ms-1], consistent with results from previous Titan model intercomparison studies (Lora et al., 2019).
-In the case of zonal winds, the upper atmosphere on Titan is dominated by eastward (superrotating) flow, driven by the upward and equatorward transport of angular momentum by the mean flow and eddies, respectively (Flasar & Achterberg, 2009; Lewis et al., 2023; Lombardo & Lora, 2023).
-Near the surface, frictional drag keeps the flow closer to Titan’s solid body rotation.
-The net gain of angular momentum from the surface at low latitudes necessitates mean westward (retrograde) winds near the surface, a feature consistently reproduced across all four simulations in our study and in prior Titan GCM studies (e.g., Lora et al., 2019).
-Although differences among simulation cases are minimal near the surface, substantial variations emerge in the stratosphere.
-In Cases 1, 2, and 4, maximum wind speeds reach approximately 150 [ms-1], whereas in Case 3 they peak at only 60 [ms-1] (not shown).
-This suggests that differences in global surface temperature structure—driven by varying thermal inertia—can influence wind patterns even at high altitudes.
-While this implies that surface-induced thermal structures may propagate upward and modulate stratospheric circulation, a detailed analysis of the vertical momentum transfer and dynamical response is beyond the scope of this study.
-For meridional winds, the flow near the surface is predominantly southward, transitioning to northward at altitudes around 2 [km].
+Across all cases, the simulations reproduce the observed wind structure in the lowest few kilometers, successfully capturing the key features of both the zonal and meridional components.
+The simulated near-surface wind magnitudes are on the order of – ms-1, consistent with results from previous Titan model intercomparison studies (Lora et al., 2019).
+While the simulated near-surface winds are centered near zero on average, the model does occasionally reproduce eastward winds consistent with observations; however, the strongest observed eastward values lie outside the model variability.
+This residual bias may reflect, at least in part, the omission of convective vertical momentum transport, though it could also indicate that the Huygens observation was atypical.
+As shown in Charnay et al.
+(2015), equatorial tropical methane storms during equinox can transport eastward winds and their associated momentum from higher altitudes downward, potentially contributing to the eastward orientation of the dune fields (Lorenz et al., 2006; Lorenz & Radebaugh, 2009).
+Titan’s middle atmosphere exhibits eastward superrotation driven by upward and equatorward angular momentum transport, while friction keeps near-surface flow close to solid-body rotation; the resulting low-latitude momentum budget favors weak mean westward winds near the surface, which our simulations reproduce and which agree with prior Titan GCM studies (Flasar & Achterberg, 2009; Lewis et al., 2023; Lombardo & Lora, 2023; Lora et al., 2019).
+Above 4 km in the troposphere, however, our simulated zonal winds are systematically weaker than observed.
+This bias likely reflects differences in the simulated thermal structure: by thermal-wind balance, understated zonal temperature gradients yield weaker shear.
+In our dry runs, the zonal surface-temperature gradients peak toward the poles (Fig.
+2), whereas including moist processes leads to more realistic gradients and winds (Lora et al., 2019).
+Although near-surface differences among Cases 1–4 are small, the stratospheric response varies substantially (not shown): peak speeds reach in Cases 1, 2, and 4 but only in Case 3, suggesting that global surface-temperature structure—set here by thermal inertia—can imprint on the circulation aloft.
+A full analysis of the vertical momentum pathways responsible for this sensitivity is left for future work, while our focus here is the lowest few kilometers where boundary-layer processes dominate and model–data agreement is generally good.
+For meridional winds, the flow near the surface is predominantly southward, transitioning to northward at altitudes around 2 km.
 This vertical wind reversal suggests the presence of a cross-equatorial Hadley circulation, characterized by rising motion in the southern summer hemisphere and descending motion in the northern winter hemisphere.
@@
 In this section, we present the same analysis of surface temperature and PBL structure as in the previous sections, but for the planned Dragonfly landing site near Selk Crater (E, N) during southern midsummer () (Barnes et al., 2021; Lorenz et al., 2021).
-For Case 4, the dunes terrain type, with a thermal inertia of [TIU], is assumed at the surface (MacKenzie et al., 2019).
+For Case 4, the dunes terrain type, with a thermal inertia of TIU, is assumed at the surface (MacKenzie et al., 2019).
 Figure 8 shows the diurnal energy balance components at the Dragonfly landing site.
 Maximum insolation is reduced compared to the Huygens landing site, as the Dragonfly site experiences northern winter conditions while the Huygens site corresponds to southern summer.
-This results in a lower mean surface temperature at the Dragonfly site, reduced by approximately 0.1 [K] relative to the Huygens site.
-The amplitude of diurnal surface temperature variations remains comparable to that at the Huygens site.
-For Case 4, the simulated diurnal amplitude is [K].
+This results in a lower mean surface temperature at the Dragonfly site, reduced by approximately 0.1 K relative to the Huygens site.
+Nevertheless, the amplitude of diurnal surface temperature variations remains comparable to that at the Huygens site.
+For Case 4, the simulated diurnal amplitude is K.
 Figure 9 shows the diurnal PBL structures at the Dragonfly landing site.
-Due to the reduced surface energy input and lower temperatures during this winter season—compared to the midsummer conditions at the Huygens landing site—the diagnosed PBL depths are shallower, reaching a maximum of approximately 500 [m], in contrast to 1,000 [m] at the Huygens site.
-Nevertheless, the overall temperature structures remain similar to those observed at the Huygens landing site.
+Due to the reduced surface energy input and lower temperatures during this winter season—compared to the midsummer conditions at the Huygens landing site—the diagnosed PBL is shallower, with maxima of m, m, m, and m across the four cases, in contrast to 1,100-1,400 m at the Huygens site.
+At night, the baseline minimum PBL depths are m, m, m, and m for Cases 1–4, respectively.
+This results in much-reduced diurnal amplitudes (minimum to maximum) of 300–400 m, compared with 800–1,100 m at the Huygens site.
 Figure 10 shows the seasonal evolution of PBL depth at the Dragonfly landing site.
-As with the Huygens landing site, higher thermal inertia surfaces lead to elevated PBL depths.
+As at the Huygens landing site, higher thermal inertia surfaces produce deeper PBLs, with maxima of , , , and m and minima of , , , and m for Cases 1–4, respectively.
 However, a key difference is observed in the seasonal symmetry.
-At the Dragonfly site, the PBL depth exhibits a more symmetric seasonal pattern: maxima occur near the equinoxes ( and ), and minima near the solstices ( and ).
+At the Dragonfly site, the PBL depth exhibits a more symmetric seasonal pattern, owing to its near-equatorial location compared to that of Huygens: maxima occur near the equinoxes ( and ), and minima near the solstices ( and ).
 In contrast, the Huygens site displays an asymmetric pattern: a pronounced minimum at the southern winter solstice (), with a higher but still suppressed depth at the southern summer solstice (), relative to the equinoctial maxima.
@@
 By combining theoretical analysis with dry GCM simulations, we identify distinct physical mechanisms that govern surface temperature variations at different timescales: subsurface heat conduction dominates the damping of diurnal variations, while atmospheric damping plays a more prominent role at seasonal timescales.
-At diurnal timescales, surface temperature amplitudes are highly sensitive to thermal inertia.
+Surface temperature amplitudes are highly sensitive to thermal inertia at diurnal timescales.
 Low thermal inertia surfaces respond inefficiently to subsurface heat conduction, leading to larger temperature variations and stronger sensible heat fluxes.
@@
 This is because the product , which characterizes the impact of subsurface heat conduction, is small compared to the atmospheric damping coefficient except under globally high thermal inertia.
-Consequently, seasonal surface temperature variations are mainly regulated by direct surface–atmosphere energy exchange through sensible heat flux, rather than by subsurface thermal diffusion.
-PBL behavior differs markedly between diurnal and seasonal timescales.
-At diurnal timescales, the PBL exhibits a regular cycle of growth and decay—rising from approximately 500 [m] in the morning to nearly 1,000 [m] in the afternoon—driven by daily variations in surface heating and cooling.
-At seasonal timescales, the PBL responds to broader shifts in insolation.
-As thermal inertia increases, the contrast between minimum and maximum surface temperature diminishes, leading the seasonal PBL to extend to a maximum depth of 2,000–3,000 [m], consistent with previous dry GCM simulations (Charnay & Lebonnois, 2012), dune spacing analysis (Lorenz et al., 2010), and the dry adiabatic lapse rate extending to 3.5 [km] observed by Voyager IRIS (Lindal et al., 1983).
-These results highlight that PBL depth is governed not only by transient energy inputs but also by the long-term thermal state of the surface and overlying atmosphere.
-Simulated winds on Titan exhibit minimal diurnal variability, which is probably linked with the small diurnal temperature swings ( [K]) compared to Earth.
-However, the GCM accurately captures seasonal wind structures, including the key zonal and meridional features observed by the Huygens probe.
+Consequently, seasonal surface temperature variations are mainly regulated by direct surface–atmosphere energy exchange through sensible heat flux, rather than by subsurface conduction.
+Thermal inertia controls on the PBL are clearly evident in the simulations.
+At diurnal timescales, the PBL undergoes a regular cycle of growth and decay—rising from a few hundred meters during the nighttime to over 1,000 meters in the afternoon—driven by daily variations in surface heating and cooling.
+Surfaces with lower thermal inertia exhibit larger diurnal fluctuations in PBL depth, accompanied by stronger surface temperature variability.
+Notably, despite stronger daytime sensible heat fluxes, lower thermal inertia surfaces have shallower daytime PBL depths.
+At night, the PBL is consistently shallower over lower thermal inertia surfaces due to rapid surface cooling.
+These features are all linked to nighttime temperature inversions, which inhibit the development of deeper boundary layers.
+At seasonal timescales, the PBL responds to broader variations in insolation.
+As with the diurnal cycle, PBL depth increases with thermal inertia, reaching maximum depths of over 2,000 m in the highest thermal inertia case.
+This is consistent with previous dry GCM simulations (Charnay & Lebonnois, 2012), dune spacing analyses (Lorenz et al., 2010), and the dry adiabatic lapse rate inferred from Voyager IRIS observations (Lindal et al., 1983).
+The influence of high thermal inertia on PBL structure is also evident in simulations incorporating a global thermal inertia map, where enhanced PBL depths are found over regions such as Xanadu.
+Simulated winds on Titan exhibit minimal diurnal variability, which is probably linked with the small diurnal temperature swings compared to Earth.
+The GCM accurately captures seasonal wind structures in the lowest few kilometers, including the key zonal and meridional features observed by the Huygens probe.
 The simulations support the interpretation by Tokano (2007) that Titan’s wind profiles are shaped primarily by global-scale circulation patterns, not Ekman dynamics as on Earth.
@@
 This transition is characteristic of Titan’s cross-equatorial Hadley circulation.
-At the Dragonfly landing site, simulations during northern midwinter indicate a shallower diurnal PBL (reaching 500 [m]) due to reduced insolation compared to the Huygens site’s midsummer conditions.
+At the Dragonfly landing site, simulations during northern midwinter show a shallower diurnal PBL—ranging from about 200 to 600 meters—due to reduced insolation compared to the midsummer conditions at the Huygens site.
 Nonetheless, temperature structures remain similar, and large-scale wind patterns are preserved, reflecting the shared equatorial latitude and the dominance of global atmospheric dynamics.
 These results suggest that Dragonfly will likely experience steady boundary layer conditions and modest winds, comparable to those encountered by Huygens.
+A previous study estimated the maximum winds (i.e., the worst-case scenario that Dragonfly might encounter) associated with dry convective vortices using the equation , obtaining a value of 2.7 ms-1 with K (Lorenz, 2021).
+This estimate was based on a simplified energy balance approach in which the atmospheric temperature, , is held constant, similar to that of Williams et al.
+(2012).
+In our simulations, which explicitly account for the full energy balance, applying Eq.
+(7) with ms-1 as in Lorenz (2021) yields (Wm-2K-1).
+Combined with a sensible heat flux ranging from 0.5 to 1.0 Wm-2 as shown in Figure 8, this results in to K and consequently to ms-1, which is in good agreement with Lorenz (2021).
 While our findings offer valuable insights, a key limitation of this study is the exclusion of methane moisture processes, which are likely to play a significant role in Titan’s climate system.
-Observations from the Huygens probe indicated that the surface was damp at the time of landing (Zarnecki et al., 2005; Lorenz et al., 2006), and our dry simulations overestimate surface temperatures by approximately 2 [K].
+Observations from the Huygens probe indicated that the surface was damp at the time of landing (Zarnecki et al., 2005; Lorenz et al., 2006), and our dry simulations overestimate surface temperatures by approximately 2 K.
 Furthermore, in several simulation cases, seasonal maxima in surface temperature and PBL depth occur at high latitudes, contrary to observational evidence indicating that Titan’s warmest regions are consistently located within 30° of the equator, as shown by Voyager IRIS data (Flasar et al., 1981; Flasar, 1998) and 13 years of Cassini CIRS observations (Jennings et al., 2016, 2019).
-Previous studies have demonstrated that the inclusion of moist processes substantially reduces the pole-to-equator surface temperature contrast, bringing simulated climate patterns into closer agreement with observations (Lora et al., 2019).
-This discrepancy highlights the importance of including latent heat processes from methane evaporation and condensation.
+Previous studies demonstrate that the inclusion of moist processes leads to improved latitudinal profiles of surface temperature and more realistic zonal wind structures, thereby bringing simulated climate patterns into closer agreement with observations (Lora et al., 2015; Mitchell & Lora, 2016; Turtle et al., 2018; Lora et al., 2019).
+This highlights the importance of including latent heat processes from methane evaporation and condensation.
 Prior studies have shown these to play a critical role in regulating Titan’s surface energy balance and atmospheric structure (Griffith et al., 2008; Mitchell et al., 2009; Lora & Ádámkovics, 2017).
@@
 (1) will be essential for future modeling efforts.
-Moreover, Huygens measurements of nearly uniform methane mole fractions in the lowest 4–5 [km] (Niemann et al., 2005, 2010; Gautier et al., 2024) suggest vigorous vertical mixing, possibly linked to the same large-scale circulation that shapes meridional wind reversals.
+Moreover, Huygens measurements of nearly uniform methane mole fractions in the lowest 4–5 km (Niemann et al., 2005, 2010; Gautier et al., 2024) suggest vigorous vertical mixing, possibly linked to the same large-scale circulation that shapes meridional wind reversals.
 All of this further motivates the inclusion of humidity and moist effects in future Titan PBL simulations.
 In summary, this study establishes a physically grounded framework for understanding Titan’s surface temperature and boundary layer variability across diurnal and seasonal timescales.
-By integrating theoretical energy balance analysis with dry GCM simulations, we demonstrate the important roles of subsurface conduction and atmospheric damping in modulating surface conditions.
+By integrating theoretical energy balance analysis with dry GCM simulations, we demonstrate the important roles of subsurface conduction and atmospheric damping in modulating surface conditions and PBL structures.
 While future work must incorporate methane-related processes to capture Titan’s full climate complexity, our results provide critical constraints for interpreting Huygens-era observations and offer quantitative projections for environmental conditions at the Dragonfly landing site, supporting mission planning and future data analysis.
@@
 We consider a solution of the form:
-where represents the complex amplitude of the temperature variation. The partial derivative of with respect to time is:
-Substituting this into (A1) gives:
-Assuming is constant (homogeneous subsurface), the equation simplifies to:
-Rearranging it, we find:
+where represents the complex amplitude of the temperature variation. Substituting this into (A1) and assuming is constant (homogeneous subsurface) gives:
 where .
@@
 Using , we find:
-At , this becomes:
-Applying the boundary condition (A2):
-Rearranging gives:
+At , and using the boundary condition (A2), we obtain
 Noting that , the subsurface temperature at any time and depth can be expressed as:
@@
 The amplitude of surface temperature oscillations is given by:
-where [Wm-2] denotes the amplitude of fluctuations in . This expression is notably independent of the specific value of the deep boundary condition , ensuring its general applicability.
+where (Wm-2) denotes the amplitude of fluctuations in . This expression is notably independent of the specific value of the deep boundary condition , ensuring its general applicability.